# HS-FPN: High Frequency and Spatial Perception FPN for Tiny Object Detection

### v1

@@ 1 Introduction
Tiny Object Detection (TOD), a subtask of general object detection, focuses on detecting tiny-sized objects that AI-TOD (Wang et al. 2021) defines as being less than $\times 16$ 16 pixels. TOD plays an important role in scenarios such as traffic sign detection, scene monitoring, unmanned aerial vehicle analysis, pedestrian detection, autonomous driving, and maritime rescue.
@@
Figure 1: Target region after filtering out low-frequency components across different ranges (top), 3D surface representations (down), and the SCR of the target (marked in red text). (a) The original image containing a tiny ship target; (b) The target neighborhood image without filtering; (c - d) Filtering out low-frequency components in the top-left corner of the DCT results, from the $\times 0$ 0 to $\times 100$ 100 region. It shows that the SCR first increases and then decreases as the filtering range is expanded.
@@
Tiny objects lack network attention. The feature responses of tiny objects are relatively weak and susceptible to interference, requiring extra attention. However, FPN applies the same processing method to features at each layer: $\times 1$ 1 convolutions to reduce the number of channels, to-down features fusion and $\times 3$ 3 convolutions to integrate outputs, without providing special treatment for tiny objects features.
@@
where $\mu_{t}$ is the average gray value of the target region (a local image region of $\times 40$ 40 pixels centered around the target in this work), and $\mu_{b}$ and $\sigma_{b}$ are the average and standard deviation of the gray value in the target’s neighborhood, respectively. Figure 1 shows the target regions, their 3D surface representations, and the SCR values after filtering out low-frequency components across different ranges. The results indicate that moderate filtering of low-frequency components can significantly improve SCR, making tiny objects more prominent.
@@ 3 Method
In this section, we detail the HS-FPN framework and its components. As shown in Figure 2, the structure of HS-FPN is similar to FPN, collecting four feature maps from the backbone network and reducing their channel dimensions to 256 using $\times 1$ 1 convolutions. These reduced-channel features are named as {${C_{2},C_{3},C_{4},C_{5}}$} which have strides of {4, 8, 16, 32} pixels relative to the input image. The feature pyramid {${P_{2},P_{3},P_{4},P_{5}}$} is generated by the top-down pathway in HS-FPN. Each lateral connection in HS-FPN contains two modules: high frequency perception module (HFP) and spatial dependency perception (SDP) module. HFP generates high frequency responses to enrich the information of tiny objects in $C_{i}$ while SDP takes {$C_{i},P_{i+1}$} as inputs using pixel-level cross attention mechanism to learn the relationships between pixels in {$C_{i},P_{i+1}$} and then using valuable spatial dependencies to enrich the features of tiny targets. Please note that all laterals of HS-FPN contain the HFP module, while only {${P_{2},P_{3},P_{4}}$} layers contain the SDP module. Finally, the output features of HS-FPN at each layer are obtained from {${P_{2},P_{3},P_{4},P_{5}}$} through separate $\times 3$ 3 convolutions for subsequent detection tasks.
@@ 3.1 High Frequency Perception Module
To address the issues mentioned in Section 1 that tiny objects have limited usable features and FPN does not pay extra attention to tiny objects, we propose the high frequency perception module (HFP). As shown in Figure 3, HFP consists of a high-frequency feature generator, a channel path (CP), and a spatial path (SP). Assuming the input feature $C_{i}$ has a size of $R^{(C\times{H_{i}}\times{W_{i}})}$, HFP first applies a high-pass filter to $C_{i}$ to obtain the high frequency responses $F_{i}$ which has the same size of $C_{i}$. Then $F_{i}$ is used respectively as input for CP and SP to generate channel attention weights $u^{CP}\in R^{(C\times{1}\times{1})}$ and spatial attention weights $u^{SP}\in R^{(1\times{H_{i}}\times{W_{i}})}$. These weights are broadcast along the spatial and channel dimensions, respectively, and element-wise multiplied with $C_{i}$. Finally, the results from CP and SP are added and then passed through a $\times 3$ 3 convolution to produce the output of HFP, denoted as ${C_{i}}^{HFP}$.
@@
To more effectively extract information from each channel, we chose to compress the spatial dimensions of $F_{i}$, ensuring that each channel of $F_{i}$ is representative. Global Average Pooling (GAP) can integrate the spatial information of the feature map to obtain an overall representation, while Global Max Pooling (GMP) captures the maximum activation value in each channel. Therefore, we use both GAP and GMP to process $F_{i}$, integrating its feature information to facilitate the subsequent calculation of each channel’s importance to tiny objects. Firstly, as shown in Figure 3, we perform a GAP and a GMP on $F_{i}$ separately, resulting in $F_{i}^{GAP}$ and $F_{i}^{GMP}$. It is important to note that we directly pool $F_{i}$ into $R^{(C\times{k}\times{k})}$ instead of pooling it to the same size of $u^{CP}$, as that results in a significant loss of information. In this paper, $k$ is generally set to 16, as when k is too small, key information is lost in both GAP and GMP; when k is too large, the effectiveness of both operations diminishes. Next, we apply the ReLU function to both $F_{i}^{GAP}$ and $F_{i}^{GMP}$ to retain positive values, and then sum across each channel to generate two one-dimensional feature vectors. These vectors are then passed through separate $\times 1$ 1 group convolutions to generate different channel scores. These scores are then concatenated and passed through another $\times 1$ 1 group convolution, producing the final channel attention weights $u^{CP}$.
Spatial Path. The SP of HFP is designed to help the model better focus on tiny object features within the spatial range, achieving an effect similar to self-attention (Vaswani et al. 2017). However, self-attention relies on similarity calculations between feature pixels to emphasize important information. Given that tiny object features are weak and FPN’s shallow feature maps contain noise, self-attention may struggle to effectively highlight tiny object features and could even amplify noise and background responses. With the help of the high-pass filter, the low-frequency homogeneous background in $F_{i}$ is filtered out and tiny object features are enhanced. Therefore, we use $F_{i}$ as a spatial mask. Specifically, a $\times 1$ 1 convolution aggregates the channel-wise information of $F_{i}$ to generate the spatial attention mask $u^{SP}$.
@@ 3.2 Spatial Dependency Perception Module
As shown in Figure 6, we first upsample $P_{i+1}$ to match the size of $C_{i}$, and name the upsampled feature $P^{u}_{i+1}$. Subsequently, we compute the Query (Q), Key (K), and Value (V) matrices through three separate $\times 1$ 1 convolutions, where Q is derived from $C_{i}$ and K, V are derived from $P^{u}_{i+1}$. Next, we divide the {Q, K, V} matrices into multiple feature blocks. For simplicity, let’s assume the dimensions of {Q, K, V} are $R^{(C\times{H_{i}}\times{W_{i}})}$. Since the detection framework doesn’t constrain input image size, the feature block size must be adjusted based on the actual input image to avoid issues with non-divisible dimensions. Please note that the input feature maps in the feature pyramid {${C_{2},C_{3},C_{4},C_{5}}$} are multiples of each other in size, so making the feature block have the same size with $C_{5}$ could ensure a perfect division. Based on this assumption, we partition {Q, K, V} into feature blocks and reshape them into $R^{(n\times({H_{5}}\times{W_{5})\times{C}})}$, where $n=\frac{H_{i}}{H_{5}}\times\frac{W_{i}}{W_{5}}$, $C$ is the number of channels in $C_{i}$, ${H_{5}}\times{W_{5}}$ is the resolution of $C_{5}$. It can be explained as: $C_{i}$ can be divided into $n$ feature blocks with each block containing ${H_{5}}\times{W_{5}}$ feature points, and each point is represented as a $C$ dimensional column vector. Then we extract corresponding feature blocks $q_{j}$ and $k_{j}$ ($1\leq{j}\leq{n},q_{j},k_{j}\in R^{((H_{5}\times{W_{5}})\times{C})}$) from {Q, K} and calculate their similarity and obtain the similarity matrix $a_{j}\in R^{((H_{5}\times{W_{5}})\times{(H_{5}\times{W_{5}})})}$. Finally, each feature point in the corresponding block $v_{j}$ from V is weighted by $a_{j}$ to generate feature blocks rich in spatial dependencies and high-level semantics.
@@ 4.1 Datasets and Metrics
[…]
@@ 4.2 Experiments Settings
Figure 7: Visualization of the feature at the $P_{2}$ level before and after added HFP and SDP to FPN separately.
@@ 4.3 Ablation Study
The effect of Spatial Dependency Perception Module. We implement SDP on each lateral layer of FPN independently to evaluate its performance, resulting in an improvement of $AP$ by 1.1 points. When combined with HFP, the $AP$ improved by 3.4 points over the original FPN, 1.2 points over HFP alone, and 2.3 points over SDP alone, showing the benefits of their combination. As shown in Figure 7 (c), compared to the input feature, the tiny object features output by SDP are more abundant.
@@ 4.5 Computational Costs
Table 5 demonstrates the computation increase of each component of HS-FPN within the entire TOD framework. The entire detection model is Cascade R-CNN with ResNet50 as the backbone and an input size of (800, 800). The Params and FLOPs for HFP mainly come from the $\times 3$ 3 convolutions in Figure 2, which results in similar computational costs for CP and SP. Given the context of the entire detection framework, HS-FPN slightly increases computational cost but significantly improves performance compared to FPN.
@@ A.3 Datasets and Metrics
[…]

### v2

@@ 1 Introduction
Tiny Object Detection (TOD), a subtask of general object detection, focuses on detecting tiny-sized objects that AI-TOD (Wang et al. 2021) defines as being less than $16\times 16$ pixels. TOD plays an important role in scenarios such as traffic sign detection, scene monitoring, unmanned aerial vehicle analysis, pedestrian detection, autonomous driving, and maritime rescue.
@@
Figure 1: Target region after filtering out low-frequency components across different ranges (top), 3D surface representations (down), and the SCR of the target (marked in red text). (a) The original image containing a tiny ship target; (b) The target neighborhood image without filtering; (c - d) Filtering out low-frequency components in the top-left corner of the DCT results, from the $0\times 0$ to $100\times 100$ region. It shows that the SCR first increases and then decreases as the filtering range is expanded.
@@
Tiny objects lack network attention. The feature responses of tiny objects are relatively weak and susceptible to interference, requiring extra attention. However, FPN applies the same processing method to features at each layer: $1\times 1$ convolutions to reduce the number of channels, to-down features fusion and $3\times 3$ convolutions to integrate outputs, without providing special treatment for tiny objects features.
@@
where $\mu_{t}$ is the average gray value of the target region (a local image region of $40\times 40$ pixels centered around the target in this work), and $\mu_{b}$ and $\sigma_{b}$ are the average and standard deviation of the gray value in the target’s neighborhood, respectively. Figure 1 shows the target regions, their 3D surface representations, and the SCR values after filtering out low-frequency components across different ranges. The results indicate that moderate filtering of low-frequency components can significantly improve SCR, making tiny objects more prominent.
@@ 3 Method
In this section, we detail the HS-FPN framework and its components. As shown in Figure 2, the structure of HS-FPN is similar to FPN, collecting four feature maps from the backbone network and reducing their channel dimensions to 256 using $1\times 1$ convolutions. These reduced-channel features are named as {${C_{2},C_{3},C_{4},C_{5}}$} which have strides of {4, 8, 16, 32} pixels relative to the input image. The feature pyramid {${P_{2},P_{3},P_{4},P_{5}}$} is generated by the top-down pathway in HS-FPN. Each lateral connection in HS-FPN contains two modules: high frequency perception module (HFP) and spatial dependency perception (SDP) module. HFP generates high frequency responses to enrich the information of tiny objects in $C_{i}$ while SDP takes {$C_{i},P_{i+1}$} as inputs using pixel-level cross attention mechanism to learn the relationships between pixels in {$C_{i},P_{i+1}$} and then using valuable spatial dependencies to enrich the features of tiny targets. Please note that all laterals of HS-FPN contain the HFP module, while only {${P_{2},P_{3},P_{4}}$} layers contain the SDP module. Finally, the output features of HS-FPN at each layer are obtained from {${P_{2},P_{3},P_{4},P_{5}}$} through separate $3\times 3$ convolutions for subsequent detection tasks.
@@ 3.1 High Frequency Perception Module
To address the issues mentioned in Section 1 that tiny objects have limited usable features and FPN does not pay extra attention to tiny objects, we propose the high frequency perception module (HFP). As shown in Figure 3, HFP consists of a high-frequency feature generator, a channel path (CP), and a spatial path (SP). Assuming the input feature $C_{i}$ has a size of $R^{(C\times{H_{i}}\times{W_{i}})}$, HFP first applies a high-pass filter to $C_{i}$ to obtain the high frequency responses $F_{i}$ which has the same size of $C_{i}$. Then $F_{i}$ is used respectively as input for CP and SP to generate channel attention weights $u^{CP}\in R^{(C\times{1}\times{1})}$ and spatial attention weights $u^{SP}\in R^{(1\times{H_{i}}\times{W_{i}})}$. These weights are broadcast along the spatial and channel dimensions, respectively, and element-wise multiplied with $C_{i}$. Finally, the results from CP and SP are added and then passed through a $3\times 3$ convolution to produce the output of HFP, denoted as ${C_{i}}^{HFP}$.
@@
To more effectively extract information from each channel, we chose to compress the spatial dimensions of $F_{i}$, ensuring that each channel of $F_{i}$ is representative. Global Average Pooling (GAP) can integrate the spatial information of the feature map to obtain an overall representation, while Global Max Pooling (GMP) captures the maximum activation value in each channel. Therefore, we use both GAP and GMP to process $F_{i}$, integrating its feature information to facilitate the subsequent calculation of each channel’s importance to tiny objects. Firstly, as shown in Figure 3, we perform a GAP and a GMP on $F_{i}$ separately, resulting in $F_{i}^{GAP}$ and $F_{i}^{GMP}$. It is important to note that we directly pool $F_{i}$ into $R^{(C\times{k}\times{k})}$ instead of pooling it to the same size of $u^{CP}$, as that results in a significant loss of information. In this paper, $k$ is generally set to 16, as when k is too small, key information is lost in both GAP and GMP; when k is too large, the effectiveness of both operations diminishes. Next, we apply the ReLU function to both $F_{i}^{GAP}$ and $F_{i}^{GMP}$ to retain positive values, and then sum across each channel to generate two one-dimensional feature vectors. These vectors are then passed through separate $1\times 1$ group convolutions to generate different channel scores. These scores are then concatenated and passed through another $1\times 1$ group convolution, producing the final channel attention weights $u^{CP}$.
Spatial Path. The SP of HFP is designed to help the model better focus on tiny object features within the spatial range, achieving an effect similar to self-attention (Vaswani et al. 2017). However, self-attention relies on similarity calculations between feature pixels to emphasize important information. Given that tiny object features are weak and FPN’s shallow feature maps contain noise, self-attention may struggle to effectively highlight tiny object features and could even amplify noise and background responses. With the help of the high-pass filter, the low-frequency homogeneous background in $F_{i}$ is filtered out and tiny object features are enhanced. Therefore, we use $F_{i}$ as a spatial mask. Specifically, a $1\times 1$ convolution aggregates the channel-wise information of $F_{i}$ to generate the spatial attention mask $u^{SP}$.
@@ 3.2 Spatial Dependency Perception Module
As shown in Figure 6, we first upsample $P_{i+1}$ to match the size of $C_{i}$, and name the upsampled feature $P^{u}_{i+1}$. Subsequently, we compute the Query (Q), Key (K), and Value (V) matrices through three separate $1\times 1$ convolutions, where Q is derived from $C_{i}$ and K, V are derived from $P^{u}_{i+1}$. Next, we divide the {Q, K, V} matrices into multiple feature blocks. For simplicity, let’s assume the dimensions of {Q, K, V} are $R^{(C\times{H_{i}}\times{W_{i}})}$. Since the detection framework doesn’t constrain input image size, the feature block size must be adjusted based on the actual input image to avoid issues with non-divisible dimensions. Please note that the input feature maps in the feature pyramid {${C_{2},C_{3},C_{4},C_{5}}$} are multiples of each other in size, so making the feature block have the same size with $C_{5}$ could ensure a perfect division. Based on this assumption, we partition {Q, K, V} into feature blocks and reshape them into $R^{(n\times({H_{5}}\times{W_{5})\times{C}})}$, where $n=\frac{H_{i}}{H_{5}}\times\frac{W_{i}}{W_{5}}$, $C$ is the number of channels in $C_{i}$, ${H_{5}}\times{W_{5}}$ is the resolution of $C_{5}$. It can be explained as: $C_{i}$ can be divided into $n$ feature blocks with each block containing ${H_{5}}\times{W_{5}}$ feature points, and each point is represented as a $C$ dimensional column vector. Then we extract corresponding feature blocks $q_{j}$ and $k_{j}$ ($1\leq{j}\leq{n},q_{j},k_{j}\in R^{((H_{5}\times{W_{5}})\times{C})}$) from {Q, K} and calculate their similarity and obtain the similarity matrix $a_{j}\in R^{((H_{5}\times{W_{5}})\times{(H_{5}\times{W_{5}})})}$. Finally, each feature point in the corresponding block $v_{j}$ from V is weighted by $a_{j}$ to generate feature blocks rich in spatial dependencies and high-level semantics.
@@ 4.1 Datasets and Metrics
Experiments are performed on two TOD datasets. The main experiments are conducted on the challenging AI-TOD (Wang et al. 2021), which provides a total of 700,621 instances in 8 categories in 28,036 aerial images. The average absolute instance size of AI-TOD is 12.8 pixels. Furthermore, we build a subset consisting of 10 categories named $DOTA_{mini10}$ from DOTA (Xia et al. 2018) and test our method on it. This subset contains 410,305 instances, with $17.3\%$ of the instances smaller than 16 pixels and $38.5\%$ smaller than 32 pixels. All metrics used in the experiments follow the AI-TOD benchmark.
@@ 4.3 Ablation Study
The effect of Spatial Dependency Perception Module. We implement SDP on each lateral layer of FPN independently to evaluate its performance, resulting in an improvement of $AP$ by 1.1 points. When combined with HFP, the $AP$ increased by 3.4 points compared to FPN, 1.2 points over HFP alone, and 2.3 points over SDP alone, showing the benefits of their combination. As shown in Figure 7 (c), the combination of HFP and SDP effectively enriches and enhances the features of tiny objects, while also suppressing some of the high-frequency noise amplified by HFP.
Figure 7: Visualization of the feature at the $P_{2}$ level after the gradual addition of HFP and SDP in FPN.
@@ 4.5 Computational Costs
Table 5 demonstrates the computation increase of each component of HS-FPN within the entire TOD framework. The entire detection model is Cascade R-CNN with ResNet50 as the backbone and an input size of (800, 800). The Params and FLOPs for HFP mainly come from the $3\times 3$ convolutions in Figure 2, which results in similar computational costs for CP and SP. Given the context of the entire detection framework, HS-FPN slightly increases computational cost but significantly improves performance compared to FPN.
@@ 5 Conclusion
Extended version. Additional motivation and details are provided in the extended version [^1] https://arxiv.org/abs/2412.10116.
@@ A.3 Datasets and Metrics
Building details. Firstly, We selected 10 balanced categories from the DOTA train set and validation set, including plane (PL), small-vehicle (SV), large-vehicle (LV), ship (SH), tennis-court (TC), baseball-diamond (BD), swimming-pool (SV), basketball-court (BC), storage-tank (ST), helicopter (HE). Second, we cleaned and organized the annotations of the original DOTA dataset. We then scaled the images in the dataset by a factor of 0.5 to 1.0, and applied parameters with a stride of 768, an overlap of 256 to split the images, with sub-images of size $1024\times 1024$ pixels. Finally, we performed HSV color space transformation and probability flipping (p-flip) on the objects of the three categories: BC, BD, HE, in order to augment the number of instances. We named the dataset constructed above as $DOTA_{mini10}$, which consists of 10,653 images and 328,086 object instances in the train set, and 2,733 images and 82,219 object instances in the validation set. Notably, $17.3\%$ of instances are smaller than 16 pixels, and $38.5\%$ are smaller than 32 pixels in $DOTA_{mini10}$, respectively.